# The taggedPBC: Annotating a massive parallel corpus for crosslinguistic investigations

### v1

@@ 3.1.1 Validation of tags via automated POS-taggers
State of the art pos-taggers have been developed for high-resource languages via the SpaCy [^25] https://spacy.io/ and Trankit [^26] http://nlp.uoregon.edu/trankit text processing libraries. These libraries have supervised pos-taggers for 16 and 31 languages respectively (excluding languages with non-roman scripts). For this comparison, each of the trained taggers was used to tag the reduced PBC verses for the respective language and then compared with the same verses tagged via the word alignment method. Correspondence with verb and noun tags between the word alignment method and these pos-taggers was on average above 75% for both SpaCy (Table 2; nouns: 0.84; verbs: 0.76) and Trankit (Table 3; nouns: 0.84; verbs: 0.77).
ISO 639-3 Language Nouns Verbs
@@
ISO 639-3 Language Nouns Verbs
@@
This comparison informs us that the translation method is relatively robust at transferring tags from English to other languages. At least for tagging Nouns and Verbs, it aligns pretty well with automatic taggers individually developed for specific languages.
@@ 3.1.2 Validation of tags via gold standard corpora
The most comparable dataset with tagged corpus data, as noted above, is the UDT. As of the v2.14 release, there were 160 languages represented in this dataset. Of this data, 106 languages contain more than 200 hand-tagged sentences, and 96 of these languages are also present in the taggedPBC. Given that the UDT contains a mixture of different genres of texts for each language, this provides a good validation of the IBM2 pos-tagging methodology outlined above. If the method used to tag the PBC results in word/tag pairs that exactly match with words and tags in the UDT for a given language, we can surmise that other words in the PBC with the same tags are highly likely to belong to the same class.
Somewhat surprisingly, we find this indeed to be the case. Despite the PBC being a rather different genre of texts than those that make up the UDT corpora, there are words with the exact same form and tag for every one of the 96 languages shared by the two corpora where the UDT has at least 200 sentences. In fact, there are only two languages with only one or fewer predicates that match in form and POS tag: Urdu (ISO 639-3: urd) which shares one verbal form (kin) between the two tagged corpora, and Apurinã (ISO639-3: apu) which shares none. When we investigate these more closely, it turns out that in the Urdu UDT corpus there are no diacritics to mark vowels, whereas the Urdu PBC corpus contains diacritics that allow for insertion of vowels in the romanized script. For Apurinã, it seems that the two corpora use different characters for different vowel sounds (so PBC popĩkari corresponds to UDT pupỹkary). Despite these differences in spelling conventions, the Urdu UDT additionally shares two nouns with the PBC (shak, din) while the Apurinã UDT shares five nouns with the PBC (tirikapi, arẽka, kãkiti, itixi, mãkatxi).
The two points of comparison (between automated taggers on the one hand and hand-tagged corpora on the other) indicate that the method used for tagging parts of speech for low-resource languages in the PBC is a valid approach, given that it largely aligns with POS taggers for high-resource languages and at the same time identifies actual arguments and predicates as found in hand-tagged data. However, “the proof is in the pudding”: does the data in the taggedPBC lend itself to actual correlations with observed properties of language? To assess this, we now turn to observation of word order that can be extracted from this dataset.
@@ 3.2 Testing features extracted from the dataset: word order
By “word order” most researchers refer to the relative order of referential arguments (nouns) versus predicates (verbs), such that it is possible to talk about “verb-initial” vs “verb-final” or “verb-medial” languages, as well as grammatical constructs like Subject and Object in relation to the position of the Verb. In order to identify these relations in corpora I use the following procedure for each language in the taggedPBC corpus:
@@
    Make lists of all the verses in the corpus, filtering out verses/sentences which do not contain both a noun and a verb.
@@
    In the resulting set, verses/sentences with a noun/argument occurring first are considered to be “noun first”, while those with a verb/predicate occurring first are considered “verb first”.
@@
To test the validity of this measure I compared the derived values with known expert judgements of word order (SV, VS, or “free”) in three typological databases (Figure 1): the World Atlas of Language Structures (WALS; Dryer & Haspelmath 2013), [^27] https://wals.info/ Grambank (Skirgård et al., 2023), [^28] https://grambank.clld.org/ and the AUTOTYP Database (Bickel et al., 2023), [^29] https://github.com/autotyp/autotyp-data. For each of these databases, I conducted a one-way ANOVA with the N1 ratio as the dependent variable, and the database category as the grouping variable for all languages shared between the PBC and the respective database (WALS: 566 languages; Grambank: 791 languages; AUTOTYP: 180 languages). [^30] Code to reproduce these results can be found at the linked repository. For each comparison, the N1 ratio significantly differentiated between SV and VS languages (WALS mean diff. 1.296, p < 0.001; Grambank mean diff. 1.19, p < 0.001; AUTOTYP mean diff. 1.386, p < 0.001), and for WALS and Grambank it additionally differentiated between SV, VS and free word orders, indicating that this ratio can be used as a proxy for word order, at least for the PBC.
@@
As an additional verification step of the N1 ratio as a proxy for word order, I trained a machine learning classifier to predict word order classifications for languages in the PBC for which word order was unknown. Since the taggedPBC data is normally distributed (see Figure 2 and additional statistical tests at the linked repository), I trained a Gaussian Naive Bayes (GNB) classifier on the 976 languages with known word order (with the N1 ratio as the feature), and then predicted on the remaining 621 languages with unknown word order. This resulted in 510 of the languages being classified as "SV", and 111 being classified as "VS". Initial tests with other classifiers gave similar results, though some (like the Decision Tree Classifier) identified up to 10 previously unclassified languages as "free". When literature searches were conducted for these languages (see Table 4), existing research identified all of the languages as having "VS" basic word order, validating the GNB classification.
@@ 4 Conclusions and future work
The preceding sections have presented a new, massive dataset of parallel text tagged for parts of speech in 1,597 languages, derived from the PBC (Mayer & Cysouw 2014) via cross-lingual POS-tag transfer. The dataset is available for research on Github as the taggedPBC, allowing for expansion/collaboration, and has been validated via comparison with taggers for high-resource languages as well as hand-tagged data from the UDT. Additionally, I have shown that the N1 ratio, a feature extracted from this dataset, differentiates between the word order (SV, VS, free) of languages, allowing for the classification of previously unclassified languages, suggesting exciting new directions for research in computational typology.
@@ 4.1 Potential for expansion
The availability of the dataset via Github means that there are many possibilities for expansion and refinement of the data. The current release provides a baseline of POS tags via crosslingual transfer using the IBM2 model, but other alignment and transfer methods could be applied to increase accuracy for individual languages or groups of languages. Additionally, manually tagging portions of the individual corpora could allow for the training of individual models on a per-language basis that would improve accuracy. In each case, it would be important to validate such approaches, while balancing accuracy with computational efficiency.
Individual datasets could also be expanded, either through tagging of additional portions of the Bible present in the larger PBC, or by adding data from other (hand-tagged) corpora. While this would affect the overall distribution of data, it might also benefit crosslinguistic investigations in various ways. Multiple levels of annotation could also be supported, either through automatic annotation via lexical databases or manual annotation, provided that these were also validated/verified in a principled manner.
@@ 4.2 Potential research directions
This dataset provides an important source of data for the growing field of computational typology. As mentioned in the footnote above, this field has primarily been focused on statistical methods whereby different features of language (or of environment) are correlated with determinations or classifications in typological databases (see Jäger 2025). This kind of investigation is limited by the availability of information on such features or classifications for particular languages, and there may be some debate regarding how to precisely categorize a language (with respect to word order, see the large number of “UNK” [unknown] classifications for word order in the Grambank database).
The dataset presented here, along with the example of extracting the N1 ratio and observing its relation to word order, provides a potential new direction for computational typology, whereby features can be directly extracted from corpora. As with the example above, datasets like the taggedPBC allow researchers to identify corpus-based features that correlate with (and/or differentiate between) expert classifications in typological databases. These features can then be used to classify previously unclassified languages, greatly expanding the coverage of such databases, and directly addressing the issue of linguistic diversity, as highlighted by Benítez-Burraco (2025) and others. Further, such corpus-derived measures allow for differing approaches to linguistic questions, such as the gradient-based word order approach proposed by Levshina et al. (2023). While the example above shows that languages do not overlap very much in terms of the N1 ratio with regard to basic word order classifications, the point remains that no language is entirely SV or VS, supporting alternative methods to answer related questions.
Some questions that this dataset is uniquely able to support investigations in are questions related to word types crosslinguistically, whether with regard to length, order, or other patterns. The fact that such questions can be queried in relation to corpora additionally allows for known effects of frequency on language (Behrens & Pfänder 2016; Bentz et al. 2017; Bybee 2006; Diessel 2016; Ellis 2016; Haspelmath 2021; Koplenig et al. 2022) to be accounted for. Other questions that could be asked, related to dependencies and phonological properties, may require additional annotation, whether automated, semi-automated, or manual. While improvements can definitely be made to the overall annotations, the taggedPBC dataset provides a solid baseline to push computational typology forward as a means of investigating universal properties of human language.

### v2

@@ 3.1.1 Validation of tags via automated POS-taggers
POS-taggers have been developed for high-resource languages, and the current state of the art is represented by the SpaCy [^26] https://spacy.io/ and Trankit [^27] http://nlp.uoregon.edu/trankit text processing libraries. These libraries have supervised POS-taggers for 16 and 31 languages respectively (excluding languages with non-roman scripts). For this comparison, each of the trained taggers in these libraries was used to tag the reduced PBC verses for the respective language and then the specific tags for arguments (NOUN, PROPN, PRON) and predicates (VERB, AUX) were compared with the same verses tagged via the word alignment method. [^28] I should note here, first, that both libraries use a variation of the universal tag set adopted by the UDT project and my methodology, as outlined in footnote 25. While there are some differences, they apply to less common word classes and do not affect arguments and predicates, which were the focus of this comparison. Second, these taggers were not developed using Biblical texts, which reduces the overall likelihood of correspondence, and makes the actual result more striking. Specifically, each word-alignment tagged verse was stripped of tags and re-tagged by the respective libraries. Then the two sets of tagged verses (between the library and the word-alignment method) were compared and the number of matching arguments/predicates was divided by the total number of arguments/predicates identified by the respective library in a given sentence. This was then averaged across all verses to provide an agreement percentage between the translation method and the respective library for the argument and predicate tags. [^29] See code at the linked repository for a more detailed breakdown of correspondence between individual POS tags in the argument and predicate classes. Correspondence with argument and predicate tags between the word alignment method and these pos-taggers was on average above 75% for both SpaCy (Table 2; arguments: 0.84; predicates: 0.76) and Trankit (Table 3; arguments: 0.84; predicates: 0.77).
ISO 639-3 Language Arguments Predicates
@@
ISO 639-3 Language Arguments Predicates
@@
This comparison informs us that the translation method agrees with existing SOTA POS-taggers more than 75% of the time, which suggests that this methodology is relatively robust at transferring tags from English to other languages. At least for tagging major word classes, it aligns pretty well with automatic taggers individually developed for specific languages.
@@ 3.1.2 Validation of tags via gold standard corpora
The most comparable dataset with tagged corpus data, as noted above, is the UDT. As of the v2.14 release, there were 160 languages represented in this dataset. Of this data, 106 languages contain more than 200 hand-tagged sentences, and 86 of these languages are also present in the taggedPBC. Given that the UDT contains a mixture of different genres of texts for each language, this provides a good validation of the IBM2 pos-tagging methodology outlined above. If the method used to tag the PBC results in word/tag pairs that exactly match with both words and tags in the UDT for a given language (i.e. for both datasets, in Yakut [ISO 639-3: sah] kiliebi is tagged as ‘NOUN’ and kel is tagged as ‘VERB’), we can surmise that other words in the PBC with the same tags are highly likely to belong to the same class.
Despite the PBC being a rather different genre of texts than those that make up the UDT corpora, this is indeed the case: there are words with the exact same form and tag for every one of the 86 languages shared by the two corpora where the UDT has at least 200 sentences. On average, each pair of corpora shares 137 nouns and 89 verbs as exact matches. In fact, there is only one language with one or fewer verbs that match in form and POS tag: Urdu (ISO 639-3: urd) which shares one verbal form (kin) between the two tagged corpora. When we investigate this more closely, it turns out that in the Urdu UDT corpus there are no diacritics to mark vowels, whereas the Urdu PBC corpus contains diacritics that allow for insertion of vowels in the romanized script. The exact matches and code to identify these correspondences can be found at the linked repository.
The two points of comparison (between automated taggers on the one hand and hand-tagged corpora on the other) indicate that the method used for tagging parts of speech for low-resource languages in the PBC is a valid approach, given that it largely aligns with POS taggers for high-resource languages and at the same time identifies actual nouns and verbs as found in hand-tagged data. However, “the proof is in the pudding”: does the data in the taggedPBC lend itself to actual correlations with observed properties of language? To assess this, we now turn to observations of word order that can be extracted from this dataset.
@@ 3.2 Testing features extracted from the dataset: word order
By “word order” most researchers refer to the relative order of referential arguments (nouns, proper nouns, pronouns) versus predicates (verbs, auxiliaries), such that it is possible to talk about “verb-initial” vs “verb-final” or “verb-medial” languages, as well as grammatical constructs like Subject and Object in relation to the position of the Verb. In the current investigation I focus on intransitive word order, which is concerned with a predicate that requires a single core argument, resulting in three possible word orders: “verb final” (SV), “verb initial” (VS), and “free” (no dominant order). In order to identify these relations in corpora I use the following procedure for each language in the taggedPBC corpus:
@@
    Make lists of all the verses in the corpus, filtering out verses/sentences which do not contain both an argument and a predicate.
@@
    In the resulting set, verses/sentences with an argument occurring first are considered to be “noun first”, while those with a predicate occurring first are considered “verb first”.
@@
To test the validity of this measure I compared the derived values with known expert judgements of intransitive word order (SV, VS, or “free”) in three typological databases (Figure 1): the World Atlas of Language Structures (WALS; Dryer & Haspelmath 2013), Grambank (Skirgård et al., 2023), and the AUTOTYP Database (Bickel et al., 2023). For each of these databases, I conducted a one-way ANOVA with the N1 ratio as the dependent variable, and the database category as the grouping variable for all languages shared between the PBC and the respective database (WALS: 646 languages; Grambank: 937 languages; AUTOTYP: 203 languages). [^30] Code to reproduce these results can be found at the linked repository. For each comparison, the N1 ratio significantly differentiated between SV and VS languages (WALS mean diff. 1.443, p < 0.001; Grambank mean diff. 1.285, p < 0.001; AUTOTYP mean diff. 1.544, p < 0.001), and for WALS and Grambank it additionally differentiated between SV/VS and free word orders, [^31] For AUTOTYP the N1 ratio significantly differentiated between SV and free word orders (p ¡ 0.001) but was marginal for VS and free languages (p = 0.054). indicating that this ratio can be used as a proxy for word order, at least for the PBC.
@@
As an additional verification step of the N1 ratio as a proxy for word order, I trained a machine learning classifier to predict word order classifications for languages in the PBC for which word order was unknown. Since the taggedPBC data is normally distributed (see Figure 2 and additional statistical tests at the linked repository), I trained a Gaussian Naive Bayes (GNB) classifier on the 1,152 languages with known word order (with the N1 ratio as the feature), and then predicted on the remaining 790 languages with unknown word order. This resulted in 655 of the languages being classified as “SV”, and 135 being classified as “VS”. Initial tests with other classifiers gave similar results, though some (like the Decision Tree Classifier) identified up to 10 previously unclassified languages as “free”. When literature searches were conducted for these languages (see Table 4), existing research identified all of the languages as having “VS” basic word order, validating the GNB classification.
@@ 4 Conclusions and future work
The preceding sections have presented a new, massive dataset of parallel text tagged for parts of speech in 1,942 languages via cross-lingual POS-tag transfer, derived partially from the PBC (Mayer & Cysouw 2014) with additional data sourced from newly-accessible Bible translations. The dataset is available for research on Github as the taggedPBC, allowing for expansion/collaboration, and the tags for nouns and verbs have been validated via comparison with taggers for high-resource languages as well as hand-tagged data from the UDT. Additionally, I have shown that the N1 ratio, a feature extracted from this dataset on the basis of arguments and predicates, differentiates between the intransitive word order (SV, VS, free) of languages, allowing for the classification of previously unclassified languages, and suggesting exciting new directions for research in computational typology.
@@ 4.1 Potential for expansion
The availability of the dataset via Github means that there are many possibilities for expansion and refinement of the data. The current release provides a baseline of POS tags via crosslingual transfer using the IBM2 model, but other alignment and transfer methods could be applied to increase accuracy for individual languages or groups of languages. As noted, there are concerns with the applicability of universal POS tags for specific languages and language families. To address this, manually tagging portions of the individual corpora could allow for the training of individual models on a per-language basis that would improve accuracy. Some guidance along these lines is presented at the linked repository to facilitate collaboration. In each case, it would be important to validate such approaches, while balancing accuracy with computational efficiency.
Individual datasets could also be expanded, either through tagging of additional portions of the Bible present in the original PBC, or by adding data from other (hand-tagged) corpora. While this would affect the overall distribution of data, it might also benefit crosslinguistic investigations in various ways. Multiple levels of annotation could also be supported, either through automatic annotation via lexical databases or manual annotation, provided that these are also validated/verified in a principled manner.
@@ 4.2 Potential research directions
This dataset provides an important source of data for the growing field of computational typology. As mentioned in footnote 9 above, this field has primarily been focused on statistical methods whereby different features of language (or of environment) are correlated with determinations or classifications in typological databases (see Jäger 2025). This kind of investigation is limited by the availability of information on such features or classifications for particular languages, and there may be some debate regarding how to precisely categorize a language (with respect to word order, see the large number of “UNK” [unknown] classifications for word order in the Grambank database).
The dataset presented here, along with the example of extracting the N1 ratio and observing its relation to word order, provides a potential new direction for computational typology, whereby features can be directly extracted from corpora. As with the example above, datasets like the taggedPBC allow researchers to identify corpus-based features that correlate with (and/or differentiate between) expert classifications in typological databases. These features can then be used to classify previously unclassified languages, greatly expanding the coverage of such databases, and directly addressing the issue of linguistic diversity, as highlighted by Ben´ıtez-Burraco (2025) and others. Further, such corpus-derived measures allow for differing approaches to linguistic questions, such as the gradient-based word order approach proposed by Levshina (2023). While the example above shows that languages do not overlap very much in terms of the N1 ratio with regard to basic intransitive word order classifications, the point remains that no language is entirely SV or VS, supporting alternative methods to answer related questions.
Some questions that this dataset is uniquely able to support investigations in are those related to word types crosslinguistically, whether with regard to length, order, or other patterns. The fact that such questions can be queried in relation to corpora additionally allows for known effects of frequency on language (Behrens & Pfänder 2016; Bentz et al. 2017; Bybee 2006; Haspelmath 2021; Koplenig et al. 2022) to be accounted for. Other questions that could be asked, related to dependencies and phonological properties, may require additional annotation, whether automated, semi-automated, or manual. While improvements can definitely be made to the overall annotations, the taggedPBC dataset provides a solid baseline to push computational typology forward as a means of investigating (universal) properties of human language.